OpenStudy (anonymous):

Line A is represented by the equation given below: x + y = 4 What is most likely the equation for line B, so that the set of equations has infinitely many solutions?

OpenStudy (anonymous):

i need help!

OpenStudy (anonymous):

multiply line A by any number except 0 and it will still be the equation of the same line. You could multiply by 2 and get 2x + 2y = 8 or multiply by -5 and get -5x - 5y = - 20.

OpenStudy (anonymous):

4x + 4y = 4 4x + y = 4 2x + 2y = 8 x + y = 8

OpenStudy (anonymous):

so would it be D?

OpenStudy (anonymous):

i ment C

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you